OpenStudy (anonymous):

Graph the circle with center at (4, 0) and radius 3. Label the center and at least four points on the circle. Write the equation of the circle.

OpenStudy (lucaz):

the equation of a circle is (x-a)^2+(y-b)^2=r^2 r is the radius and (a,b) the center of the circle

OpenStudy (lucaz):

the equation is (x-4)^2+(y-0)^2=3^2 or x^2+y^2-8x=-7

OpenStudy (anonymous):

I got the first one, now what do I do after that?

OpenStudy (lucaz):

the graph is easy, you know the center and the radius, to find the four points you just pick a value of x and solve for y

OpenStudy (lucaz):

actually, it's easier using the pythagorean theorem